# Enhancing Trust in LLM-Generated Code Summaries with Calibrated Confidence Scores

### v2

@@ Abstract.
A good summary can often be very useful during program comprehension. While a brief, fluent, and relevant summary can be helpful, it does require significant human effort to produce. Often, good summaries are unavailable in software projects, thus making maintenance more difficult. There has been a considerable body of research into automated AI-based methods, using Large Language models (LLMs), to generate summaries of code; there also has been quite a bit work on ways to measure the performance of such summarization methods, with special attention paid to how closely these AI-generated summaries resemble a summary a human might have produced. Measures such as BERTScore and BLEU have been suggested and evaluated with human-subject studies.
However, LLM-produced summaries can be too long, irrelevant, etc: generally, too dissimilar to what a human might say. Given an LLM-produced code summary, how can we judge if a summary is good enough? Given some input source code, and an LLM-generated summary, existing approaches can help judge brevity, fluency and relevance; however, it’s difficult to gauge whether an LLM-produced summary sufficiently resembles what a human might produce, without a “golden” human-produced summary to compare against. We study this resemblance question as a calibration problem: given just the summary from an LLM, can we compute a confidence measure, that provides a reliable indication of whether the summary sufficiently resembles what a human would have produced in this situation? We examine this question using several LLMs, for several languages, and in several different settings. Our investigation suggests approaches to provide reliable predictions of the likelihood that an LLM-generated summary would sufficiently resemble a summary a human might write for the same code.
@@
ccs: Software and its engineeringccs: Software and its engineering Software maintenance toolsccs: Software and its engineering Documentation
@@ 1. Introduction
Summaries and comments in code help maintainers understand code. In an empirical study, Hu et al surveyed developers (Hu et al., 2022) in several major software development organizations about software summarization practices. The survey responses indicated that developers value code summaries, and, furthermore, that they would appreciate tools to automatically generate such summaries for them. Automated Code Summarization has been a very active area of research (Zhang et al., 2022; Zhu and Pan, 2019) for quite a while. Deep learning, for code and natural language, has helped advance summarization performance to new highs, with the current state-of-the-art achieved via the use of in-context learning (or prompt engineering) with large language models (LLMs) (Ahmed et al., 2024; Nashid et al., 2023).
@@
Figure 1. Example of two LLM-generated summaries with nearly identical, raw confidences but differing quality. The generated summary on the left is dissimilar to what the developer wrote, but the confidence is high. The generated summary on the right is similar to what the developer wrote, and the raw confidence is high, as it should be; however, the raw confidence is also high on the left! Our method preserves the high confidence for the good summary but produces a low confidence for the poor one.
While neural models are increasingly effective at code summarization, there is a fundamental and well-known problem with LLMs $\ldots$ they make mistakes (Jesse et al., 2023; Pan et al., 2024; Tambon et al., 2024). Studies (McBurney and McMillan, 2015; Shi et al., 2022) have found that automated code summarization methods are promising, but often fail to produce “acceptable” summaries. Thus, given a generated summary, it would be useful to have an indication as to whether the generated summary is a “acceptable” one. Summary quality is a human judgement; there have been studies of how humans evaluate code summaries, which we review in the paper. For reasons clarified below in Section 2, and supported by several prior publications (Hu et al., 2022; Stapleton et al., 2020) we use “acceptable” similarity to human-written summaries as our primary correctness measure.
In our study, we found that state-of-the-art approaches (using CodeLlama) sometimes produce Java code summaries at acceptable levels of similarity to human-written summaries (using measures that previous studies (Haque et al., 2022) have found to be reliable proxies of human similarity judgements). In our study, we find that (arguably) acceptable levels of similarity are attained between 10-30% of the time. This suggests that about 70-80% of the time, these approaches are producing summaries that humans may not find acceptably similar to human summaries. But how do we know whether the summary will be acceptable or not without the presence of the reference? Can raw model confidence, taken from model output without change, help us in such a scenario? Figure 1 provides two example generated summaries, however, one resembles the developer’s comment, and one doesn’t! Although the dissimilar summary doesn’t contain any falsehoods about the code, it still omits information the method’s developer thought necessary to include. Certainly, the similar one is more desirable than the dissimilar one. Unfortunately, the raw confidence does not show much difference (0.82 vs. 0.81)! Thus, the raw confidence is not particularly helpful to help determine which one is more acceptable.
@@
Thus, summaries provided by LLMs are sometimes good and useful, but not always. Given this situation, it is desirable for LLM-based code-summarizers to provide a more reliable indication of confidence, that a specific summary generated from code, is sufficiently similar to what a human might have produced. If this confidence measure is reliable, then a summary produced with high confidence can be accepted verbatim, and used directly to help understand the code; summaries at low confidence might be simply rejected; and summaries at a medium confidence might be considered partially useful, predicating some further examination of the code. This is where calibration (discussed in Section 2) becomes relevant. Calibration refers to the agreement between the predicted confidence and the actual outcome. A well-calibrated model will produce probability estimates that match the true likelihood of an event. For example, if a model predicts that an event has a 80% chance of occurring, then that event should indeed occur approximately 80% of the time across many predictions with the same probability. The model’s predicted confidence in Figure 1 shows that the LLMs are typically not well-calibrated for this task and are thus less reliable.
@@
In this paper, we study the issue of providing such a confidence measure, which is a reliable indicator of how likely a generated summary is correct, viz., sufficiently similar to a human-generated summary. For instance, in Figure 1, had our approach been used, then the more desirable outcome has a higher confidence (0.87 vs 0.42) $\ldots$; thus, our approach would allow for better decision making as to when a summary could be relied upon.
@@
  • We introduce the problem of calibration for code summarizers; we frame “correctness” of generated summaries (with motivation drawn from the literature) in terms of being sufficiently similar to human-generated summaries.
  • Using several thresholds of similarity measures, corresponding to several justifiable settings of “correctness” (i.e., similarity to human-produced summaries), we examine the calibration of LLMs.
  • We show how LLMs augmented with in-context learning, and using modern rescaling techniques, can provide very good calibration, for the code summarization task.
  • While later tokens in a code summary have very high probabilities and exhibit overconfidence, earlier tokens are better calibrated.
Our findings provide a way for developers to make well-justified decisions regarding the use of (sometimes inaccurate) summaries generated by LLMs.
@@ 2. Background & Motivation
Code summarization is important for code understanding and maintenance (Sridhara et al., 2010); building tools that can automatically summarize a given piece of code (method or snippet) is an important problem that has received considerable attention (Zhu and Pan, 2019; Zhang et al., 2022). as output. Neural approaches have been applied to code summarization. Initial approaches used pre-training + fine-tuning, with models such as CodeBERT (Feng et al., 2020), PLBART (Ahmad et al., 2021), PolyGlot (Ahmed and Devanbu, 2022b), and CodeT5 (Wang et al., 2021). Current approaches rely on Large Language Models (LLMs) like the GPT-3 model (e.g., Code-Davinci-002) (Brown et al., 2020; Ahmed and Devanbu, 2022a; Nashid et al., 2023); with such powerful LLMs, in-context learning techniques e.g few-shotting (Brown et al., 2020; Ahmed and Devanbu, 2022a) and semantic augmentation (Ahmed et al., 2024), attain state-of-the art code summarization performance (Ahmed and Devanbu, 2022a; Ahmed et al., 2024) without any additional training. So, in this paper, our focus is on LLM-based code summarizers that use in-context learning.
Despite the progress in code summarization, code summaries are still in general, produced by imperfect, flawed neural models. In our studies, we found that less than 30% of generated summaries could be considered high quality. In this setting, users could benefit from guidance as to whether a potentially flawed summary from an imperfectly performing neural model can be relied upon.
@@ 2.1. Calibration: How to use unreliable outputs from flawed models
Neural (more generally, Machine-learning) models are probabilistically optimized over the training set. On a distinct test set , they may not always produce high-quality results. How then, is one to use these flawed models in practice? This question relates to the concept of Calibration.
@@
Brier Score (Brier, 1950) is the mean squared difference between the predicted probabilities and the actual outcomes; it measures (on average) how well a sample predicted confidence align with the actual correctness on that sample. A perfectly calibrated model would have a Brier Score of 0, indicating that its predicted probabilities match the observed outcomes exactly. Mathematically, the Brier Score $B_{r}$ for a set of predictions, of sample size $N$ is calculated as follows:
@@
We note that Brier scores should be gauged relative to a base rate. A naïve model, which always outputs (as confidence) the empirical base rate $p$, (for example, if it rains 10% of the days, the model predicts rain with $p=0.1$ confidence every day) will achieve a score called the reference Brier score, $B_{ref}$, which analytically is $p(1-p)$. Thus, if it rains about 10% of the days, the base rate is $0.1$; a model that always outputs $0.1$ as its confidence will score $0.1*0.9=0.09$ Brier. Evidently, $B_{ref}\in(0,0.25]$.
@@
Brier Skill Score is a normalized measure $\in(-\infty,1]$ for assessing the reliability of a model’s confidence, relative to a naïve model always just providing the base rate as the confidence. Such a model reaches the $B_{ref}$ Brier score of $p(1-p)$, which can be quite low, depending on $p$. For a two-class prediction, with a 10% overall positive rate: a naïve predictor always guessing one class with 10% confidence yields $B_{ref}\approx 0.09$. A better-calibrated model can achieve a Brier Score lower than this unskilled $B_{ref}$ value; a bad model, could do worse! The Skill Score (SS) quantifies this improvement (or decline) compared to the baseline $B_{ref}$, calculated as:
@@
A positive SS (perfect score = 1.0) indicates improvement over the baseline, while a negative SS suggests predictions worse than the baseline. Even small positive values of SS can sometimes indicate good skill. The German weather forecasting service, Deutsche Wetterdienst, sets a minimum threshold of 0.05 for a Skill Score to indicate good forecast quality reference (Wet, 2024). Another reference point comes from the American data journalism site 538, which reports a skill score of approximately 0.13 in forecasting World Cup games (Wezerek, 2023)
@@
ECE. (Expected Calibration Error) (Naeini et al., 2015) is also used to measure calibration. It is somewhat more complex: it averages the difference between the predicted confidence and actual correctness rates across different confidence levels. Intuitively, on a reliability diagram (such as Figure 2) it amounts to measuring the (normalized) area between the diagonal ideal and the actual empirical bars. A lower ECE indicates better calibration, meaning the predicted probabilities closely match the actual rates. Conversely, a higher ECE indicates poorer calibration, indicating a mismatch between predicted and actual rates.
@@
To compute ECE, we first partition the dataset into $n$ bins $B_{i},i=1\ldots n$ based on confidences (predicted probabilities). For instance, the bins could range from confidence levels of [0-0.1) to [0.9-1.0). For each bin, we calculate the average confidence, $conf(B_{i})$ (predicted probability), and empirical accuracy, $acc(B_{i})$ (proportion of true positives). Following this, we determined a weighted difference by (i) computing the absolute difference between the average confidence and accuracy in each bin and (ii) weighting by the proportion of samples in each bin. Finally, we sum the weighted differences across all bins to get the ECE. ECE provides a more nuanced evaluation of calibration across different confidence levels compared to metrics like the Brier Score, which aggregates errors across all samples without considering confidence levels. This makes ECE particularly useful for understanding how well a model’s confidence estimates align with its performance at different confidence levels. Mathematically, the ECE is calculated in three steps below (Note: $B_{i}$ denotes the set of elements in each bin; $\hat{o_{i}}$ is an indicator varible, $=1$ if the prediction is right, and $=0$ if wrong; $\hat{p_{i}}$ is the confidence (probability) associated with the prediction)
@@ 2.2. Confidence in a generated summary
Thus far, our discussion of calibration has focused settings where there is a single, binary prediction, with an associated confidence (probability) measure. With generated code summaries, the situation is different. Here, we have a sequence multiple tokens (usually sub-tokens) generated from a softmax layer; from this layer, we can sequentially select a most likely token, with an associated probability. This would be “greedy” decoding; other methods are possible. Whatever method is used, one can obtain a generated code summary, which contains a sequence of tokens, each with an associated probability. Since calibration measures rely on a single confidence associated with an entire output, we would need a way to summarize the entire sequence of probabilities. The simplest approach is to just take a mean (geometric or arithmetic) of this sequence.
It’s worth noting here that LLMs are per se trained to auto-regressively re-generate the sequence of tokens in the training set, at the highest possible confidence. Thus, (if a given test sample isn’t too “distant” from the training data) it is reasonable to expect that the per-token confidence, is a plausible indication that a generated sequence is, in some sense, appropriate in context, relative to the data the model is trained on.
We discuss further details in the methods section.
@@ 2.3. Correctness of a generated summary
For a binary prediction (e.g. “Will it rain tomorrow”) the notion of correctness is simple (either it rained or didn’t). For a generated sequence of tokens constituting a code summary, what exactly is correctness? Exact match? Partial match? Same “meaning”? Fortunately, there has been a quite a bit of work on evaluations of code summaries (Roy et al., 2021; Shi et al., 2022; Haque et al., 2022; Mastropaolo et al., 2024).
@@
Hu et al (Hu et al., 2022) conducted a human-subject study to identify the criteria that affect practitioners’ judgement of a code summarization technique’s performance. The important evaluation criteria (for summaries) they identify include content adequacy (which they define roughly as semantic similarity between the input code and the given summary), concisenes, fluency, useful information content not in code, and similarity to original human-written comments. It should be noted that of the above 5 criteria, the first 4 are properties (functions) of the input code and the candidate summary, and potentially could be evaluated with just those two; the last requires access to a human-written summary, which is typically not available when one is using a LLM to generate code summaries. Moreover, other studies have reinforced the importance of similarity to human-written summaries. An additional human-subject study, by Stapleton et al (Stapleton et al., 2020) compared how human-written and machine-written summaries helped with comprehension tasks. They found that human-written summaries were associated with significantly better comprehension performance. This body of human-subject studies suggests that generated summaries that better resemble human-written summaries are more desirable. In this paper, we therefore focus on similarity to human-written summary as our “goodness” criteria.
@@
Even prior to He et al’s and Stapleton et al’s human-subject studies, researchers have sought the right metrics (Haque et al., 2022; Roy et al., 2021; Shi et al., 2022; Stapleton et al., 2020) to measure the similarity of a machine-generated summary to a human-produced summary. Metrics such as BLEU (Papineni et al., 2002), ROGUE (Lin, 2004), METEOR (Banerjee and Lavie, 2005) have been used to measure lexical (word n-gram level) similarity; but these have been criticized as not being well-aligned to human judgements of similarity in the case of code summaries (Stapleton et al., 2020). Haque et al. (Haque et al., 2022) empirically compared several different lexical & semantic metrics of similarity, and found that certain embedding-based metrics such as BERTScore are better-correlated with human evaluations of summary similarity (to a reference) than purely lexical measures. They report that SentenceBERT has the highest correlation with human evaluations of summary similarity.
@@
We re-analyzed all the data & metrics used by Haque et al (Haque et al., 2022), from a somewhat different perspective: Can we reliably predict sufficient similarity, of a model-generated summary? For our perspective we found that both BERTScore and SentenceBERT could be useful. Note that in our setting, we seek to predict sufficiently high semantic similarity (above a threshold value, as discussed in  subsection 4.4) of a generated summary, without knowing the Gold (human-generated) summary. We would like to generate a well-calibrated confidence signal (associated with the generated summary) that is a reliable indicator of the empirical likelihood of sufficient semantic similarity between generated summary and the gold summary (were it available) being high enough. Our re-analysis of the Haque et al data is discussed in methods section, § 4.4.
@@
An Example Scenario: To illustrate our approach: consider a user $\mathcal{U}$ who is dealing with code $C$ that has no human-written summary; $\mathcal{U}$ then generates a summary $S$ from $C$, using a neural model. How would $\mathcal{U}$ evaluate $S$, relative to the code $C$? Considering Hu et al’s 5 criteria: $\mathcal{U}$ could judge fluency and conciseness just from reading $S$. Recently, Mastropaolo et al. (Mastropaolo et al., 2024) introduced SIDE, a way to automatically measure content adequacy of a candidate summary using just the code and the candidate summary (and no human-written reference). Finally, regarding useful content not in the code, $\mathcal{U}$ could read $C$ and $S$, and make a qualitative judgement as to whether $S$ contains relevant information not obvious from $C$. Finally, turning to similarity to developer-written comments, $\mathcal{U}$ currently has no way develop a reliable level of confidence in $S$, without access to the developer-written summary. Given the importance prior work places on similarity to human-written summary, we focus our attention on developing a well-calibrated confidence predictive of such similarity. Now, provided with a well-calibrated confidence for $S$ , $\mathcal{U}$ could make a decision that best suits their needs. Thus, if $C$ is critical for $\mathcal{U}$’s work, and is really important to understand, $\mathcal{U}$ might accept only very high-confidence summaries, and ask a human in all other cases. In cases where $C$ is less critical, and a surface understanding is sufficient, a moderate level of confidence in $S$ might be sufficient. If $S$ is delivered at a very low confidence, $\mathcal{U}$ might just reject it.
@@
The above scenario illustrates the value of associating a well-calibrated confidence with a summary generated by a neural model. A different, but analogous scenario, regarding the use of well-calibrated confidences by a user considering whether to use model-generated code, has been discussed in prior work (Huang et al., 2023; Johnson et al., 2023; Spiess et al., 2025). We are now ready to present our research questions.

### v3

# Calibration of Large Language Models on Code Summarization

@@ Abstract.
A brief, fluent, and relevant summary can be helpful during program comprehension; however, such a summary does require significant human effort to produce. Often, good summaries are unavailable in software projects, which makes maintenance more difficult. There has been a considerable body of research into automated AI-based methods, using Large Language models (LLMs), to generate summaries of code; there also has been quite a bit of work on ways to measure the performance of such summarization methods, with special attention paid to how closely these AI-generated summaries resemble a summary a human might have produced. Measures such as BERTScore and BLEU have been suggested and evaluated with human-subject studies.
However, LLM-generated summaries can be inaccurate, incomplete, etc : generally, too dissimilar to one that a good developer might write. Given an LLM-generated code summary, how can a user rationally judge if a summary is sufficiently good and reliable? Given just some input source code, and an LLM-generated summary, existing approaches can help judge brevity, fluency and relevance of the summary; however, it’s difficult to gauge whether an LLM-generated summary sufficiently resembles what a human might produce, without a “golden” human-produced summary to compare against. We study this resemblance question as a calibration problem: given just the code & the summary from an LLM, can we compute a confidence measure, that provides a reliable indication of whether the summary sufficiently resembles what a human would have produced in this situation? We examine this question using several LLMs, for several languages, and in several different settings. Our investigation suggests approaches to provide reliable predictions of the likelihood that an LLM-generated summary would sufficiently resemble a summary a human might write for the same code.
@@
copyright: ccdoi: 10.1145/3729400journalyear: 2025journal: PACMSEjournalvolume: 2journalnumber: FSEarticle: FSE130publicationmonth: 7conference: ACM on Software Engineering, Volume 2, Number FSE; June 23–27, 2025; Trondheim, Norwayisbn: 2994-970X/2025/7ccs: Software and its engineeringccs: Software and its engineering Software maintenance toolsccs: Software and its engineering Documentation
@@ 1. Introduction
Code summaries and comments help maintainers understand code. (Hu et al., 2022) surveyed developers in several major software development organizations on software summarization practices. The survey responses indicated that developers value code summaries and, furthermore, that they would appreciate tools to automatically generate such summaries for them. Automated Code Summarization has been a very active area of research (Zhang et al., 2022; Zhu and Pan, 2019) for quite a while. Deep learning has helped advance summarization performance to new heights, with the current state-of-the-art achieved via the use of in-context learning (or prompt engineering) with large language models (LLM) (Ahmed et al., 2024; Nashid et al., 2023).
@@
Figure 1. Example of two LLM-generated summaries with nearly identical confidences but differing quality. The generated summary on the left is dissimilar to what the developer wrote, but the confidence is high. The generated summary on the right is similar to what the developer wrote, and the confidence is high, as it should be; however, the confidence is also high on the left! Our method preserves the high confidence for the good summary but produces a low confidence for the poor one.
Although LLMs are increasingly effective at code summarization, sometimes even surpassing the quality of human-written summaries (discussed in Section 6), they still have lapses in quality (Su and McMillan, 2024; Afrin et al., 2025). For example, 46.3% of the time, humans judge LLM-generated code summaries as worse than developer-written ones, at explaining the role of a method (Su et al., 2024). In our study, 70-80% of the time, LLMs produce summaries humans may not find acceptably similar to human summaries. As LLMs generate code summaries at scale, inaccurate or incomplete code summaries even 1% of the time may significantly hurt program comprehension if relied upon.
Thus, given an AI-generated summary, it would be helpful to have an indication as to whether that summary is an “acceptable” one. For reasons clarified below in Section 2, and due to the importance that prior work places on it (Hu et al., 2022; Shi et al., 2022; Mastropaolo et al., 2024), we use “acceptable” similarity to human-written summaries as our primary correctness measure. When used together other metrics, this measure helps holistically evaluate a code summary’s quality.
But how do we gauge whether the summary will be acceptable or not without the presence of the reference? In NLP (Guo et al., 2017), model log-probabilities have been widely studied as confidence measures for classification tasks, with various scaling methods proposed to improve their reliability. More recently, similar techniques have been applied to assess LLM calibration in SE tasks (Huang et al., 2025), including program repair (Zhou et al., 2024; Zhang et al., 2023; Spiess et al., 2025). Can the raw model confidence, directly derived from the model’s output, reliably indicate whether the summary will be acceptable, without access to a reference?
@@
If this confidence measure is reliable, then a summary produced with high confidence can be accepted verbatim and used directly to help understand the code; summaries at low confidence might be simply rejected; and summaries at a medium confidence might be considered partially useful, requiring some further examination of the code. This is where calibration (discussed in Section 2) becomes relevant. Calibration refers to the agreement between the predicted confidence of an event and the actual event outcome. A well-calibrated model will produce probability estimates that match the true likelihood of an event. For example, if a model predicts that an event has a 80% chance of occurring, then that event should indeed occur approximately 80% of the time across many predictions with the same probability. The model’s high confidence in a low-quality summary in Figure 1 illustrates how LLMs can be poorly-calibrated for code summarization.
@@
In this paper, we study the issue of providing a confidence measure, that is a reliable indicator of how likely a generated summary is correct, viz., sufficiently similar to a human-written summary. For instance, in Figure 1, had our approach been used, then the more desirable outcome in fact comes with a higher confidence (0.87 vs 0.42); thus, our approach would support better decision-making regarding the reliability of a summary.
@@
  • We introduce the problem of calibration for code summarizers; we frame “correctness” of generated summaries in terms of being sufficiently similar to human-generated summaries.
  • Using several thresholds of similarity measures, corresponding to several settings of “correctness”, we examine the calibration of LLMs.
  • We show how LLMs augmented with in-context learning, and using modern rescaling techniques, can provide better calibration for the code summarization task.
  • While later tokens in a long code summary have very high probabilities and exhibit overconfidence, earlier tokens are better calibrated.
Our findings provide a way for developers to make well-justified decisions regarding the use of (sometimes low-quality) summaries generated by LLMs.
@@ 2. Background & Motivation
Code summarization is important for code understanding and maintenance (Sridhara et al., 2010); building tools that can automatically summarize a given piece of code (method or snippet) is an important challenge that has received considerable attention (Zhu and Pan, 2019; Zhang et al., 2022). Neural approaches have been applied to code summarization. Earlier approaches used pre-training + fine-tuning, with models such as CodeBERT (Feng et al., 2020), PLBART (Ahmad et al., 2021), PolyGlot (Ahmed and Devanbu, 2022b), and CodeT5 (Wang et al., 2021). Current approaches exploit Large Language Models (LLMs)  (Brown et al., 2020; Nashid et al., 2023; Su and McMillan, 2024; Afrin et al., 2025; Su et al., 2024); with such powerful LLMs, in-context learning techniques e.g few-shotting (Brown et al., 2020; Ahmed and Devanbu, 2022a), chain-of-thought (Sun et al., 2024), and semantic augmentation (Ahmed et al., 2024), attain state-of-the art code summarization performance without any additional training. In this paper, our focus is on LLM-based code summarizers that use in-context learning and semantic augmentation. Despite the progress in code summarization, code summaries are still produced by imperfect models.
@@ 2.1. Calibration: How to use unreliable outputs from flawed models
Neural models are probabilistically optimized over the training set. On a distinct test set, they may not always produce high-quality results. How then, is one to use these flawed models in practice? This question relates to the concept of Calibration.
@@
Brier Score [^1] See (Brier, 1950) for details. is the mean squared difference between the predicted probabilities and the actual outcomes; it measures (on average) how well (given a sample) predicted confidence aligns with the actual correctness on that sample. A perfectly calibrated model would have a Brier Score of 0, indicating that its predicted probabilities match the observed outcomes exactly. Mathematically, the Brier Score $B_{r}$ for a sample set of $N$ predictions is calculated as follows:
@@
Brier scores should always be gauged relative to a base rate. A naïve model, which always outputs (as confidence) the empirical base rate $p$, (for example, if it rains 10% of the days, the model predicts rain with $p=0.1$ confidence every day) will achieve a score called the reference Brier score, $B_{ref}$, which analytically is $p(1-p)$. Thus, if it rains about 10% of the days, the base rate is $0.1$; a model that always outputs $0.1$ as its confidence will score $0.1*0.9=0.09$ Brier. Evidently, $B_{ref}\in(0,0.25]$.
@@
Brier Skill Score is a normalized measure $\in(-\infty,1]$ for assessing the reliability of a model’s confidence, relative to a naïve model always just providing the base rate as the confidence, and a skill of $B_{ref}$. A better-calibrated model will achieve a Brier Score lower than this unskilled $B_{ref}$ value; a bad model, could do worse! The Skill Score (SS) quantifies this improvement (or decline) compared to the baseline $B_{ref}$, calculated as:
@@
A positive SS indicates improvement over the baseline, while a negative SS suggests predictions worse than the baseline. A perfect skill score is 1.0; but even small positive values of SS can sometimes indicate good skill. The German weather forecasting service, Deutsche Wetterdienst, sets a minimum threshold of 0.05 for a Skill Score to indicate good forecast quality reference (Wet, 2024). Another reference point comes from the American data journalism site 538, which reports a skill score of approximately 0.13 in forecasting World Cup games (Wezerek, 2023).
@@
ECE. (Expected Calibration Error) (Naeini et al., 2015) is another, rather more complex, calibration measure. It averages the difference between the predicted confidence and actual correctness rates across different confidence levels. Intuitively, on a reliability diagram (such as Figure 2) it amounts to measuring the (normalized) area between the diagonal ideal and the actual empirical bars. A lower ECE indicates better calibration, meaning the predicted probabilities closely match the actual rates. Conversely, a higher ECE indicates poorer calibration, meaning a mismatch between predicted and actual rates.
@@
To compute ECE, we first partition the dataset into $n$ bins $B_{i},i=1\ldots n$ based on confidences (predicted probabilities). For example, the bins could range from confidence levels of [0-0.1) to [0.9-1.0). For each bin, we calculate the average confidence, $conf(B_{i})$ (predicted probability), and empirical accuracy, $acc(B_{i})$ (fraction of true positives). Following this, we determined a weighted difference by (i) computing the absolute difference between the average confidence and accuracy in each bin and (ii) weighting by the proportion of samples in each bin. Finally, we sum the weighted differences across all bins to get the ECE. ECE provides a more nuanced evaluation of calibration across different confidence levels compared to metrics like the Brier Score, which aggregates errors across all samples without considering confidence levels. This makes ECE particularly useful for understanding how well a model’s confidence estimates align with its performance at different confidence levels. Mathematically, the ECE is calculated in three steps below (Note: $B_{i}$ denotes the set of elements in each bin; $\hat{o_{i}}$ is an indicator varible, $=1$ if the prediction is right, and $=0$ if wrong; $\hat{p_{i}}$ is the confidence (probability) associated with the prediction)
@@ 2.2. Confidence in a generated summary
Thus far, our discussion of calibration has focused settings where there is a single, binary prediction, with an associated confidence (probability) measure. With generated code summaries, we have a sequence of tokens generated from a softmax layer; from this layer, we can sequentially select a most likely token, with an associated probability. This would be “greedy” decoding; other methods are possible. Whatever method is used, one obtains a generated code summary, which contains a sequence of tokens, each with an associated probability. Since calibration measures rely on a single confidence associated with an entire output, we need a way to summarize the entire sequence of probabilities. The simplest approach is to just take a mean (geometric or arithmetic) of this sequence. We discuss further details in the methods section.
@@ 2.3. Correctness of a generated summary
For a binary prediction (e.g. “Will it rain tomorrow”) the notion of correctness is simple (either it rained or didn’t). For a generated sequence of tokens constituting a code summary, what exactly is correctness? Exact match? Partial match? Similarity of “meaning”? Fortunately, there has been a quite a bit of work on evaluations of code summaries (Roy et al., 2021; Shi et al., 2022; Haque et al., 2022; Mastropaolo et al., 2024).
@@
Prior work emphasizes evaluating code summaries by their similarity to human-written summaries (Haque et al., 2022; Shi et al., 2022; Stapleton et al., 2020), in conjunction with other metrics like SIDE (Mastropaolo et al., 2024). Human-written summaries capture content actual developers deemed useful for program comprehension; similarity to them thus serves as one of multiple signals for summary quality. (Hu et al., 2022) conducted a human-subject study to identify the criteria that affect practitioners’ judgment of a code summarization technique’s performance. The important evaluation criteria include content adequacy, conciseness, fluency, useful information content not in code, and similarity to original human-written comments. It should be noted that of the above 5 criteria, the first 4 could potentially be evaluated using just the source code and generated summary; the last requires access to a human-written summary, which is typically not available when one is using a LLM to generate code summaries. In this paper, we therefore focus on similarity to human-written summary as our “goodness” criteria.
@@
Researchers have sought the right metrics to measure the similarity of a machine-generated summary to a human-produced summary  (Haque et al., 2022; Roy et al., 2021; Shi et al., 2022; Stapleton et al., 2020). Metrics such as BLEU (Papineni et al., 2002), ROGUE (Lin, 2004), METEOR (Banerjee and Lavie, 2005) have been used to measure lexical (word n-gram level) similarity; but these have been criticized as not being well-aligned to human judgements of similarity in the case of code summaries (Stapleton et al., 2020).  (Haque et al., 2022) empirically compared several different lexical & semantic metrics of similarity, and found that certain embedding-based metrics such as BERTScore are better-correlated with human evaluations of summary similarity (to a reference) than purely lexical measures. They report that SentenceBERT has the highest correlation with human evaluations of summary similarity.
@@
We re-analyzed all the data & metrics used by  (Haque et al., 2022), from a somewhat different, binary decision perspective: Can we reliably predict sufficient similarity, of a model-generated summary? For our perspective we found that both BERTScore and SentenceBERT could be useful. Note that in our setting, we seek to predict sufficiently high semantic similarity (above a threshold value, as discussed in  subsection 4.4) of a generated summary, without knowing the Gold (human-generated) summary. We would like to generate a well-calibrated confidence signal (associated with the generated summary) that is a reliable indicator of the empirical likelihood of sufficient semantic similarity between generated summary and the gold summary (were it available) being high enough. Our re-analysis of the Haque et al data is discussed in methods section, § 4.4.
@@
An Example Scenario: To illustrate our approach: consider a user $\mathcal{U}$ who is dealing with code $C$ that has no human-written summary; $\mathcal{U}$ then generates a summary $S$ from $C$, using a neural model. How would $\mathcal{U}$ evaluate $S$, relative to the code $C$? Considering Hu et al’s 5 criteria: $\mathcal{U}$ could judge fluency and conciseness just from reading $S$. Recently,  (Mastropaolo et al., 2024) introduced SIDE, a way to automatically measure the alignment of summary $S$ to $C$ using just the code and the candidate summary (and no human-written reference); it was found to strongly predict human ratings of content adequacy. Finally, turning to similarity to developer-written comments, $\mathcal{U}$ currently has no way develop a reliable level of confidence in $S$, without access to the developer-written summary. Given the importance prior work places on similarity to human-written summary, we focus our attention on developing a well-calibrated confidence predictive of such similarity. Finally, regarding useful content not in the code, $\mathcal{U}$ would have to read $C$ and $S$, or potentially use our confidence measure as a proxy; developer-written comments we compare to can incorporate working knowledge of the software process outside the code. Now, provided with a well-calibrated confidence for $S$, $\mathcal{U}$ could make a decision that best suits their needs. Thus, if $C$ is critical for $\mathcal{U}$’s work, and is really important to understand, $\mathcal{U}$ might accept only very high-confidence summaries, and ask a human in all other cases. In cases where $C$ is less critical, and a surface understanding is sufficient, a moderate level of confidence in $S$ might be sufficient. If $S$ is delivered at a very low confidence, $\mathcal{U}$ might just reject it.
@@
The above scenario illustrates the value of associating a well-calibrated confidence with a summary generated by a neural model. A different, but analogous scenario, regarding the use of well-calibrated confidences by a user considering whether to use model-generated code, has been discussed in prior work (Huang et al., 2025; Johnson et al., 2023; Spiess et al., 2025). We are now ready to present our research questions.